OpenStudy (anonymous):

Explain why: lim x-> -infinity 1 / (4^x + 1) = 1 ?

OpenStudy (anonymous):

I can simplify it down to: \[\frac{ 1 }{ \infty + 1 }\] Doesn't this equate to 0?

OpenStudy (anonymous):

Okay, for this problem we need to break it down into two parts. First... think about the Lim of 4^x We can easily see that as x gets larger, 4^x gets even larger. What's hard to visualize is when x becomes negative. For this kind of problem, I'd either graph it, or plug a few problems into my calculator. 4^0 = 1 4^(-1) = 1/4 4^(-2) = 1/8 4^(-3) = 1/16 So as x goes to negative infinity... 4^x gets really small.. in in short. Zero. So really what you have is \[\frac{ 1 }{ 0 + 1 }\] Does that help?

OpenStudy (anonymous):

The trick is when you see a negative power, think of it as it raised to the inverse power. \[4^{-x} = \frac{ 1 }{ 4^x }\]